# Combining Evidence Across Filtrations

### v3

@@ Appendix D Other types of adjusters
[…]
@@
As evident from Figure 8, when using e-processes to construct sequential tests (i.e., binary decisions), zero adjusters can provide an additional constant factor of statistical power. For this reason, we utilize the zero adjuster $\mathsf{A}_{\mathsf{zero},1}$ for the power comparison plots in Section G.1.
@@
D.2 Generalized adjusters based on spines are not $\mathfrak{e}$-lifters
@@
Appendix E A bibliographic note on terminology
@@
Further, as noted in Vovk, 2021a (, remark 1), it is a coincidence that the subject of Getoor and Sharpe, (1972)’s study was termed “conformal martingales” (unrelated to Vovk, 2021a ; Vovk, 2021b ’s “conformal test martingales” that we discuss in Example 1).
Appendix F Proofs and additional lemmas
F.1 Proof of Lemma 1
@@
$\xi_{T}=\left(\bigcup_{t\in\mathbb{N}}(\xi_{t}\cap\{T=t\}\right)\cup\left(\xi_{\infty}\cap\{T=\infty\}\right)\subseteq\bigcup_{t\geq 1}\xi_{t}.$ (44)
@@ F.2 Proof of Theorem 2
F.3 Two characterization lemmas and the proof of Theorem 3
@@
Appendix G Additional empirical results
G.1 How much power do we actually lose?
@@
We focus on the randomness testing setup of Section 4.1 and consider combining the UI e-process with the conformal test martingale. For the alternative, we consider two families. First is the Markovian family $Q^{\mathsf{Markov}}(\delta)=\mathsf{Markov}(\mu,\mu+\delta)$, where the two parameters specify the transition parameters $p_{1|0}$ and $p_{1|1}$ respectively. (The initial state is either 00 or $1$ with equal probability.) Second is the changepoint family $Q^{\mathsf{change}}(\delta)$, which repeatedly switches between IID $\mathsf{Ber}(\mu)$ and $\mathsf{Ber}(\mu+\delta)$ every 100 steps. In either case, $Q(\delta)$ is IID if and only if $\delta=0$.
@@
Our goal is to compare the e-process obtained by combining the adjust-then-combine approach, that is, $\bar{\mathfrak{e}}_{t}=\frac{1}{2}\left[\mathfrak{e}_{t}^{\mathsf{UI}}+\mathsf{A}((\mathfrak{e}_{t}^{\mathsf{conf}})^{*})\right]$, with the simple mean (without adjusters) $\bar{m}_{t}=\frac{1}{2}\left[\mathfrak{e}_{t}^{\mathsf{UI}}+\mathfrak{e}_{t}^{\mathsf{conf}}\right]$. Recalling that the simple mean is not an e-process (in any filtration), we now compare these two in terms of their mean rejection rates and rejection times. For this experiment, we only consider the rejection rate (and do not monitor the process over time); thus, we utilize the “zero” adjuster $\mathsf{A}_{\mathsf{zero},1}$, defined in Section D.1, which zeros out small evidence to achieve even higher power under the alternative.
@@
We can also plot both the statistical power of the induced test, as well as the e-power (Vovk and Wang,, 2024), against the sample size. Given an alternative distribution $Q$, we define the e-power of an e-process $(\mathfrak{e}_{t})_{t\geq 0}$ at time $t$ is defined as $\mathbb{E}_{Q}[\log\mathfrak{e}_{t}]$. It provides a different notion of statistical power focusing on how quickly the e-process can grow under the alternative.
@@
G.2 An empirical comparison with existing approaches for the $h$-step-ahead forecast evaluation problem
@@
Appendix H Cooperative skepticism: A game-theoretic protocol for combining evidence
@@
Appendix I Randomized adjustments of e-processes across filtrations
@@
I.1 The lift-then-randomize procedure
@@
[…]
@@
for $\text{for any $\mathbb{F}$-stopping time $\tau$,}\quad\tilde{p}_{\tau}^{\mathsf{ltr}}:=\frac{U}{\mathsf{A}(\mathfrak{e}_{\tau}^{*})}\wedge 1\quad\text{is a p-value for $\mathcal{P}$}.$ (60)
@@
I.2 Lifting the randomized Ville’s inequality
@@
I.3 The randomize-then-lift procedure is not $\mathbb{F}$-anytime-valid, generally
@@
[…]
@@
[…]

### v4

@@ Appendix E Other types of adjusters
where $\kappa>0$ is a hyperparameter to be chosen a priori. We refer to these adjusters as “zero” adjusters as they completely zero out smaller values of $\mathfrak{e}$. (For example, if $\kappa=1$, then $\mathsf{A}_{\mathsf{zero},1}(\mathfrak{e})=0$ for any $\mathfrak{e}<\exp(2)\approx 7.39$.) Elementary calculus shows that $\mathsf{A}_{\mathsf{zero},\kappa}(\mathfrak{e})$ is indeed an admissible adjuster, or that $\int_{1}^{\infty}\mathsf{A}_{\mathsf{zero},\kappa}(\mathfrak{e})d\mathfrak{e}=1$, for any fixed $\kappa>0$.
@@
As evident from Figure 8, when using e-processes to construct sequential tests (i.e., binary decisions), zero adjusters can provide an additional constant factor of statistical power. For this reason, we utilize the zero adjuster $\mathsf{A}_{\mathsf{zero},1}$ for the power comparison plots in Section H.1.
@@
E.2 Generalized adjusters based on spines are not $\mathfrak{e}$-lifters
@@
Appendix F A bibliographic note on terminology
@@
Further, as noted in Vovk, 2021a (, remark 1), it is a coincidence that the subject of Getoor and Sharpe, (1972)’s study was termed “conformal martingales,” which are unrelated to Vovk, 2021a ; Vovk, 2021b ’s “conformal test martingales” that we discuss in Example 1.
Appendix G Proofs and additional lemmas
G.1 Proof of Lemma 1
@@
$\xi_{T}=\left(\bigcup_{t\in\mathbb{N}}(\xi_{t}\cap\{T=t\})\right)\cup\left(\xi_{\infty}\cap\{T=\infty\}\right)\subseteq\bigcup_{t\geq 1}\xi_{t}.$ (53)
@@ G.2 Proof of Theorem 2
G.3 Two characterization lemmas and the proof of Theorem 3
@@
Appendix H Additional empirical results
H.1 How much power do we actually lose?
@@
We focus on the randomness testing setup of Section 4.1 and consider combining the UI e-process with the conformal test martingale. For the alternative, we consider two families. First is the Markovian family $Q^{\mathsf{Markov}}(\delta)=\mathsf{Markov}(\mu,\mu+\delta)$, where the two parameters specify the transition parameters $p_{0\to 1}$ and $p_{1\to 1}$ respectively. (The initial state is either 0 or $1$ with equal probability.) Second is the changepoint family $Q^{\mathsf{change}}(\delta)$, which repeatedly switches between IID $\mathsf{Ber}(\mu)$ and $\mathsf{Ber}(\mu+\delta)$ every 100 steps. In either case, $Q(\delta)$ is IID if and only if $\delta=0$.
@@
Our goal is to compare the e-process obtained by combining the adjust-then-combine approach, that is, $\bar{\mathfrak{e}}_{t}=\frac{1}{2}\left[\mathfrak{e}_{t}^{\mathsf{UI}}+\mathsf{A}((\mathfrak{e}_{t}^{\mathsf{conf}})^{*})\right]$, with the simple mean (without adjusters) $\bar{m}_{t}=\frac{1}{2}\left[\mathfrak{e}_{t}^{\mathsf{UI}}+\mathfrak{e}_{t}^{\mathsf{conf}}\right]$. Recalling that the simple mean is not an e-process (in any filtration), we now compare these two in terms of their mean rejection rates and rejection times. For this experiment, we only consider the rejection rate (and do not monitor the process over time); thus, we utilize the “zero” adjuster $\mathsf{A}_{\mathsf{zero},1}$, defined in Section E.1, which zeros out small evidence to achieve even higher power under the alternative.
@@
We can also plot both the statistical power of the induced test, as well as the e-power (Vovk and Wang,, 2024; Ramdas and Wang,, 2025), against the sample size. Given an alternative distribution $Q$, we define the e-power of an e-process $(\mathfrak{e}_{t})_{t\geq 0}$ at time $t$ is defined as $\mathbb{E}_{Q}[\log\mathfrak{e}_{t}]$. It provides a different notion of statistical power focusing on how quickly the e-process can grow under the alternative.
@@
H.2 An empirical comparison with existing approaches for the $h$-step-ahead forecast evaluation problem
@@
Appendix I Cooperative skepticism: A game-theoretic protocol for combining evidence
@@
Appendix J Randomized adjustments of e-processes across filtrations
@@
J.1 The lift-then-randomize procedure
@@
$\mathfrak{e}_{\tau}\quad\overset{\text{$\mathfrak{e}$-lifting}}{\underset{\text{(e-to-e)}}{\longrightarrow}}\quad\mathsf{A}(\mathfrak{e}_{\tau}^{*})\quad\overset{\text{$U$-rand.}}{\underset{\text{(e-to-p)}}{\longrightarrow}}\quad\frac{U}{\mathsf{A}(\mathfrak{e}_{\tau}^{*})}\wedge 1.$ (68)
@@
[…]
@@
J.2 Lifting the randomized Ville’s inequality
@@
J.3 The randomize-then-lift procedure is not $\mathbb{F}$-anytime-valid, generally
@@
$\mathfrak{e}_{\tau}\quad\overset{\text{randomized e-to-p}}{\underset{U\sim\mathsf{Unif}[0,1]}{\longrightarrow}}\quad\tilde{\mathfrak{p}}_{\tau}^{\mathsf{rtl}}:=\frac{U}{\mathfrak{e}_{\tau}}\wedge 1\quad\overset{\text{p-to-e}}{\longrightarrow}\quad\tilde{\mathfrak{e}}_{\tau}^{\mathsf{rtl}}:=\mathsf{C}\left(\frac{U}{\mathfrak{e}_{\tau}}\wedge 1\right).$ (73)
@@
$P\left(\mathfrak{e}_{\tau}\geq\frac{U}{\alpha}\right)$ $=\int_{0}^{1}P\left(\mathfrak{e}_{\tau}\geq\frac{u}{\alpha}\right)du$ (77)
$\leq\int_{0}^{1}\left(\frac{\alpha}{u}\wedge 1\right)du$ (78)
$=\int_{0}^{\alpha}du+\int_{\alpha}^{1}\frac{\alpha}{u}du=\alpha+\alpha\log(1/\alpha).$ (79)